A volleyball is spiked so that its incoming velocity of + 5.0 m/s is changed to an outgoing velocity of -15 m/s. The mass of the volleyball is 0.35 kg. What impulse does the player apply to the ball? Remember an impulse is a change in momentum.


1
Expert's answer
2021-06-08T09:35:41-0400

The information given is

  • the initial velocity is "V_{i}=+5.0\;\text{m}/\text{s}"
  • The final velocity is "V_{f}=-15.0\;\text{m}/\text{s}"
  • The mass is "m=0.35\;\text{Kg}"


The impulse is ( the change of momentum )

"I=\Delta P\\
I=P_{f}-P_{i}\\
I=m\;V_{f}-m\;V_{i}\\"


Evaluating numerically.

"I=m\;V_{f}-m\;V_{i}\\
I=0.35\;\text{Kg}\times-15.0\;\text{m}/\text{s}-0.35\;\text{Kg}\times +5.0\;\text{m}/\text{s}\\
I=-5.25\;\text{Kg}\;\text{m}/\text{s}-1.75\;\text{Kg}\;\text{m}/\text{s}\\
I=-7\;\text{Kg}\;\text{m}/\text{s}"


Answer.

The impulse applied by the player is "\displaystyle \color{red}{\boxed{I=-7\;\text{Kg}\;\text{m}/\text{s}}}"